# references.bib

### v1

@@ Abstract
This paper evaluates single-view mesh reconstruction models for creating digital twin environments in robot manipulation. Recent advances in computer vision for 3D reconstruction from single viewpoints present a potential breakthrough for efficiently creating virtual replicas of physical environments for robotics contexts. However, their suitability for physics simulations and robotics applications remains unexplored. We establish benchmarking criteria for 3D reconstruction in robotics contexts, including handling typical inputs, producing collision-free and stable reconstructions, managing occlusions, and meeting computational constraints. Our empirical evaluation using realistic robotics datasets shows that despite success on computer vision benchmarks, existing approaches fail to meet robotics-specific requirements. We quantitively examine limitations of single-view reconstruction for practical robotics implementation, in contrast to prior work that focuses on multi-view approaches. Our findings highlight critical gaps between computer vision advances and robotics needs, guiding future research at this intersection.
@@ I Introduction
Recent advances in computer vision have produced remarkable improvements in open-vocabulary 3D mesh reconstruction from single image inputs [liu_zero-1--3_2023, liu_one-2-3-45_2023, tang_lgm_2025]. These breakthroughs raise a compelling question: can such single-view reconstruction models be leveraged out-of-the-box to create accurate digital twin environments for robotic manipulation without relying on significant post-processing? Although a growing number of recent works employ such 3D reconstruction models, each has limitations in terms of reconstruction accuracy, support for physical simulation, or experimental evaluation [mu_robotwin_2024, agarwal_scenecomplete_2024, yao_cast_2025, katara_gen2sim_2024]. The ability to efficiently lift real-world observations into accurate simulations could revolutionise how robots interact with and learn from their environments. Single-view approaches are particularly valuable in robotics contexts where collecting multiple viewpoints requires costly physical movement, delays decision-making, or may be entirely infeasible due to environmental constraints or occlusions.
Creating digital twins in simulation -- virtual replicas of physical environments -- has traditionally been a labour-intensive process requiring specialist training and manually designed assets. This has posed significant hurdles to using simulators in an online fashion such as model-predictive control [qin_survey_2003, hansen_temporal_2022], planning [alterovitz_robot_2016, suomalainen_survey_2022], online safety evaluation [berkenkamp_safe_2017, o_kelly_scalable_2018], and real-time learning [smith_demonstrating_2023, bohlinger_gait_2025] for robot manipulation, as obtaining accurate state estimations for constructing digital twin environments in simulation remains non-trivial. A general, automated, and single-view real-to-sim pipeline would map the observed scene into a simulation instance by reconstructing scene objects as individual meshes. If successful, the prospects of scene-agnostic, consistent, and robust dynamics prediction from online observation through simulation are enticing, as these are data-intensive aspects of learned world models [ball_augmented_2021, prasanna_dreaming_2024]. In this work, we formulate desiderata to empirically evaluate whether state-of-the-art single-view reconstruction models are suitable for real-to-sim applications in robotics.
Real-to-sim has received significant attention with several works aiming to leverage recent reconstruction methods in robotics. However, most require dense view-points or scene scans [barcellona_dream_2024, han_re3sim_2025, li_robogsim_2024, lou_robo-gs_2024, wu_rl-gsbridge_2024, qureshi_splatsim_2024, zhu_vr-robo_2025, jia_discoverse_2024, torne_reconciling_2024, patel_real--sim--real_2024, pfaff_scalable_2025] while the practical limitations of assuming complete scene coverage have not yet been critically examined. Other works rely on only single input views [mu_robotwin_2024, agarwal_scenecomplete_2024, yao_cast_2025, katara_gen2sim_2024] but only produce domain-randomised environments instead of exact digital twins, do not perform physics simulation, or, while mentioning potential robotics applications, do not describe any robot experiments at all. We describe these works and their limitations in greater detail in Section III-G.
@@
This backdrop of real-to-sim literature motivates the question whether recent advancements in single-view mesh reconstruction models can be effectively applied to create accurate digital twin environments for online, real-time, and high-throughput robotic manipulation. To this end, we formulate desiderata that are particular to robotics and simulation for manipulation and concern the type of input data, computational resources, and overall coherence of the reconstructed scene. We empirically evaluate state-of-the-art 3D reconstruction models against these desiderata by running a series of experiments using data that is typically encountered in robotics deployment scenarios.
@@
Our aim is to guide and inform robotics practitioners and computer vision researchers about existing capability gaps, helping them select tools suited to their tasks while being aware of current limitations. Consequently, inspired by robotics, we are focusing our evaluation on reconstruction models that:
@@
Work with single-view image inputs. It is generally impractical to assume that an object can always be observed from multiple meaningful viewpoints a-priori. In some cases, such as entering a new room or opening a drawer, estimating suitable viewpoints and physically reaching them is not trivial due to space constraints and collision avoidance during planning [meng_two-stage_2017, breyer_closed-loop_2022]. In other cases, such as books in a shelf, gaining additional viewpoints is impossible without interacting with and rearranging objects in the scene [bohg_interactive_2017, pollayil_planning_2021, miao_safe_2023].
@@
Produce complete meshes for individual objects. Most robotics simulation engines operate on mesh representations [todorov_mujoco_2012, coumans_pybullet_2016, makoviychuk_isaac_2021]. While other types of 3D representation, such as point clouds or voxels, are supported in some cases, meshes remain the de-facto standard for rigid body simulation.
@@
Are trained to perform open-vocabulary reconstruction. Historically, 3D datasets have been scarce and small [chang_shapenet_2015, collins_abo_2022, wu_omniobject3d_2023]. As a result, most 3D reconstruction approaches focused on category-specific reconstruction, such as chairs or cars, to limit the task’s scope. This has changed in recent years, with ever-larger 3D datasets [deitke_objaverse-xl_2023] and 2D data being used for 3D reconstruction [wang_score_2023, poole_dreamfusion_2022]. We are particularly interested in such recent open-vocabulary reconstruction models that promise to produce accurate 3D models of highly heterogeneous object classes, as they allow state estimation in diverse robotics scenarios like households without needing to know a priori what objects are present.
Our systematic evaluation of works fitting these criteria complements existing surveys that attempt to exhaustively categorise 3D reconstruction models either generally [maxim_survey_2021, kantarci_survey_2022, bai_survey_2024, wang_deep_2024], or with a focus on specific 3D representations such as NeRFs [remondino_critical_2023], Gaussian Splatting [fei_3d_2024, zhu_3d_2024], or Meshes [chen_review_2023, lei_whats_2024]. We are not looking to provide an exhaustive overview of the 3D reconstruction model zoo but instead fundamentally address the question whether single-view reconstruction performance of existing models is sufficient to be used in robotics. [zhu_3d_2024, irshad_neural_2024] offer extensive overviews of 3D reconstruction in robotics, but mostly focus on scanning approaches that require substantial scene coverage from the input data stream. In contrast, we focus specifically on single-view reconstruction models and run experiments on real-world datasets to evaluate whether the reported performance on computer vision datasets transfers to realistic robotics domains. [zhang_text--3d_2023] have a somewhat similar motivation but focus on three text-to-3D generative models, evaluate them towards use in mobile robotics, and investigate several optimisations to improve the latency vs reconstruction quality trade-off. We instead evaluate a broad and representative range of image-to-3D models for real-to-sim reconstruction.
In Section II, we delineate desiderata for single-view 3D reconstruction models that robotics domains necessitate. While this is not a survey, we nevertheless provide a brief taxonomy of existing 3D reconstruction approaches in Section III. Even though this work does not focus on multi-view approaches, such works are included for completeness and we refer the reader to relevant surveys, where appropriate. In Section IV, we systematically evaluate a selection of single-view 3D reconstruction models towards these desiderata to see to what extent existing models are readily applicable in the robotics domain. Finally, in Section V we summarise and discuss our findings and position them in the broader context of current research directions.
@@ II Desiderata
For existing open-vocabulary 3D reconstruction models to be effectively integrated into digital twin environments that leverage physics simulators, they must fulfil several key requirements. These requirements ensure that the reconstructed models are not only visually accurate but also physically meaningful, allowing for realistic interactions within a simulation. Furthermore, deployment in the real world results in constraints regarding the available input data and available computational resources. In this section, we motivate several requirements that are key for digital twinning on robotic platforms in the wild.
@@
Robot manipulation is highly sensitive to accuracy in perceptual system components. For successful robot grasping and manipulation, surface estimates should be within 1mm of the true object surface [morgan_vision-driven_2021, zha_semantic_2020]. Some use-cases, such as assembly or part insertion, can even require sub-millimetre tolerances [ota_autonomous_2024, niu_tolerance-guided_2021, lim_grasping_2023]. For our evaluation setting of household environments, we set the target accuracy to 2mm.
These stringent accuracy requirements, however, may be challenging to meet with current 3D reconstruction models due to a potential domain gap in training data. Real-world robots are often equipped with RGB-D cameras which offer both RGB and depth perception at relatively low cost [tadic_perspectives_2022]. Yet, these cameras typically do not match the resolution and colour profile of images commonly found in 3D reconstruction training datasets [wu_omniobject3d_2023, deitke_objaverse-xl_2023]. The difference is especially noticeable when an object occupies only a small part of the image plane, as is common when observing an entire scene. Consequently, the 3D reconstruction training data distributions [tochilkin_triposr_2024, liu_syncdreamer_2023, liu_zero-1--3_2023] differ substantially from the distribution of inputs encountered during deployment in the real world, potentially resulting in less accurate reconstructions and artifacts [barcellona_dream_2024]. While some approaches like [jiang_real3d_2024] co-train on both real-world images and synthetic data to improve reconstruction accuracy, it remains an open question whether this approach can yield satisfactory results for the precision demands of robotics applications.
@@
Physics simulators often use broad- and narrow-phase collision handling to detect collisions between objects’ collision volumes. These collisions are then resolved through penalty- or constraint-based methods that modify the current force acting on the object, its velocity, or its position [todorov_mujoco_2012, coumans_pybullet_2016, makoviychuk_isaac_2021]. Many 3D reconstruction approaches process objects individually which can result in collisions when placing these reconstructions back into the scene [tochilkin_triposr_2024, liu_syncdreamer_2023, liu_zero-1--3_2023]. This is a direct result of insufficient supervision signal from the immediate vicinity of the reconstructed objects. Scenes that collide already during static scene reconstruction fall apart when performing physics steps in the simulator due to the forces and constraints that act upon colliding objects. In principle, one could selectively deactivate collisions between objects to prevent this but that could lead to further complications in downstream simulation and cannot resolve situations where collisions are necessary to preserve the integrity of a scene, such as objects resting on supporting surfaces.
Desideratum 2: No object should violate space occupied by other objects or parts of the environment. This is necessary for the reconstructed scene to be physically stable and to allow for accurate simulation of object interactions.
@@
This desideratum is based on the assumption that observed scenes are static as long as nothing interacts with them. We can consequently assume that all objects are in a stable, resting state and do not move when advancing simulated physics [ni_phyrecon_2024, guo_physically_2024, chen_atlas3d_2024, yan_phycage_2024]. Object pose and an object’s surroundings are necessary conditioning factors as otherwise the reconstruction process could ignore object and scene parts that are critical for object stability in the scene. Especially occluded objects parts and parts with little surface area tend to be ignored during reconstruction, and such inaccuracies can lead to physical instability [ni_phyrecon_2024].
Desideratum 3: The reconstructed meshes should offer physically stable poses within 5 tilt of their scene pose, meaning that all objects are in a stable resting state and do not move when advancing simulated physics. This is crucial for ensuring accurate simulation of object interactions and maintaining the integrity of the reconstructed scene.
@@
Reconstruction models must accurately resolve partial object occlusions that are ubiquitous in real-world environments [miao_safe_2023, wang_learning_2025]. Without proper occlusion handling, digital twins would contain incomplete objects, leading to incorrect simulation physics and failed manipulation attempts. Some 3D reconstruction models run a separate occlusion completion model in their pipeline which can lead to computational overhead and accumulation of reconstruction error [agarwal_scenecomplete_2024]. Other works assume that objects tend to be symmetrical or follow other manual shape priors [kong_vmap_2023, li_rico_2023]. Object stability, as described in the previous desideratum, can give valuable clues about the structure of occluded parts.
Desideratum 4: The reconstructed meshes should resolve partial occlusions with Chamfer distances no more than 10% higher than those of visible regions when compared to ground truth. This ensures that the reconstructed objects are complete and can be accurately simulated, even in challenging partially-occluded scenarios.
@@
Many 3D reconstruction models that rely on iteratively optimising a shape representation take several minutes or even hours to reconstruct objects or scenes [liu_zero-1--3_2023, qian_magic123_2023, ni_phyrecon_2024, yan_phycage_2024]. As this reconstruction pipeline needs to be run every time a new object is observed, it represents a significant computational cost when deploying (mobile) robots in online environments. Other approaches are able to produce shape reconstructions within few seconds by relying on a single forward pass [xu_instantmesh_2024, boss_sf3d_2024, tochilkin_triposr_2024, liu_one-2-3-45_2023, liu_one-2-3-45_2024]. Naturally, latency requirements differ considerably with the type of task that is to be completed. For real-time systems operating at multiple-Hertz frequency [simpkins2012real, bylow_real-time_2013, izadi_kinectfusion_2011], however, these models are likely too slow, especially since compute time accumulates if multiple objects are to be reconstructed. We define a target latency of 2s for reconstructing an entire scene consisting of multiple objects, because the evaluation data used in this paper concerns less time-critical applications, such as household environments. While one could argue that pre-computing reconstructions and retrieving them from a database when needed might address computational concerns, this approach is generally impractical in online environments like households where object positions, orientations, and even the objects themselves change frequently. Another perspective is given by a possible accuracy/efficiency trade-off, as some computationally more expensive models produce lower-error reconstructions. This would indicate that the practitioner has to make a judgement call as to what reconstruction error and computational cost can be afforded.
Additionally, robots generally have limited on-board compute due to power and payload restrictions [swaminathan_benchmarking_2024]. At the same time, many 3D reconstruction models utilise computationally expensive operations, such as diffusion [xu_instantmesh_2024, liu_syncdreamer_2023] or iterative optimisation [liu_zero-1--3_2023, wang_score_2023, qian_magic123_2023], rendering them unable to be run on-device without significantly increased latency. Running reconstruction models on dedicated servers and streaming the results to the robotic platform is often not feasible due to latency issues, network dependence, and potential security or privacy concerns [abdulsalam_security_2022].
Desideratum 5: The reconstruction pipeline should be able to produce a complete scene reconstruction within 2s, even when multiple objects are present. This ensures that the reconstruction process is efficient and can keep up with the online nature of real-world environments.
@@
While our work is not a survey, it requires us to provide an overview of relevant parts of the literature to motivate our approach and model selection, as well as to position our findings in the broader context of the 3D reconstruction literature. In this section, we present a focused taxonomy of 3D reconstruction approaches. Given the rapidly evolving nature of this research area with new methods appearing frequently, we cannot comprehensively cover all existing approaches. Instead, we focus on representative methods that illustrate key capabilities and limitations relevant to robotic manipulation requirements. This framework highlights how current approaches align with or fall short of the practical needs of robotic systems operating in real-world environments. Specifically, we look at different 3D representations that are used for reconstruction (Section III-A), various input data requirements (Section III-B), different prediction paradigms (Section III-C), different representation abstraction levels (Section III-D), the consideration of physical properties and parameters during reconstruction (Section III-E), training data properties (Section III-F), and how recent reconstruction approaches have been leveraged in a real-to-sim fashion in robotics (Section III-G). Table I gives an overview of a selection of existing reconstruction models and their properties.

### v2

# Is Single-View Mesh Reconstruction Ready for Robotics?

@@ Abstract
This paper evaluates single-view mesh reconstruction models for their potential in enabling instant digital twin creation for real-time planning and dynamics prediction using physics simulators for robotic manipulation. Recent single-view 3D reconstruction advances offer a promising avenue toward an automated real-to-sim pipeline: directly mapping a single observation of a scene into a simulation instance by reconstructing scene objects as individual, complete, and physically plausible 3D meshes. However, their suitability for physics simulations and robotics applications under immediacy, physical fidelity, and simulation readiness remains underexplored. We establish robotics-specific benchmarking criteria for 3D reconstruction, including handling typical inputs, collision-free and stable geometry, occlusions robustness, and meeting computational constraints. Our empirical evaluation using realistic robotics datasets shows that despite success on computer vision benchmarks, existing approaches fail to meet robotics-specific requirements. We quantitively examine limitations of single-view reconstruction for practical robotics implementation, in contrast to prior work that focuses on multi-view approaches. Our findings highlight critical gaps between computer vision advances and robotics needs, guiding future research at this intersection.
@@ I Introduction
Simulating future world states from perception is central to robot planning and learning [1, 2]. Recent advances in learnt world models - neural networks trained to predict forward dynamics from raw observations - have shown promise for model-based control, planning, and exploration [3, 4]. However, such models suffer from poor generalisation and opaque failure modes, especially under distribution shift, and require vast training data and offer limited physical interpretability [5, 6]. To address these issues, one might choose to bypass learnt dynamics models by instead modelling dynamics with a digital twin in simulation, a geometric reconstruction of the currently observed physical scene or object suitable for robotic simulation and interaction, enabling robots to reason about the world using physics-based simulators.
Creating such digital twins in simulation has traditionally been a labour-intensive process requiring specialist training and manually designed assets. This complexity poses substantial challenges for robotic manipulation, where the ability to rapidly create simulation environments from real-world observations is critical for applications such as real-time planning, control, and learning. These challenges have hindered the integration of simulation into online robotics workflows such as model-predictive control [7, 8], planning [9, 10], online safety evaluation [11, 12], and real-time learning [13, 14] for robot manipulation, as obtaining accurate state estimations for constructing digital twins in simulation remains non-trivial.
A general, automated scene reconstruction pipeline for robotic manipulation would directly map RGB(-D) observations of a cluttered or structured scene into a simulation instance by reconstructing scene objects as individual, physically plausible, and complete 3D meshes. If successful, the prospects of scene-agnostic and robust dynamics prediction from online observation through simulation are enticing, as these capabilities could address the fundamental data efficiency, generalisation, and robustness limitations of learned world models [5, 6].
Recent advances in real-to-sim, the process of transforming sensor observations of a physical scene - typically an RGB(-D) image - into a structured, physics-compatible simulation environment, leverage multi-view 3D reconstruction methods to create simulation-compatible environments [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. While promising, these methods inherently assume complete scene coverage from multiple viewpoints. This progress has been paralleled by emerging computer vision works that produce remarkable improvements in category-agnostic, single-view, full-shape 3D mesh reconstruction - i.e., models that infer the complete object or scene geometry, including unobserved surfaces, from single-view input [26, 27, 28]. In this setting, no feature matching cues can be exploited due to the lack of multi-view observations, and all geometric reasoning must be learnt from data during model training.
Figure 1: Illustration of a general Real2Sim2Real pipeline for robotic manipulation. A single RGB(-D) view is used for scene reconstruction in a physics simulator. The simulation instance can then be used for planning, model-predictive control, or real-time learning and the resulting motion is executed in the real world.
This backdrop raises a compelling question: can the multi-view requirement in prior works be relaxed by leveraging single-view, full-geometry reconstruction models out-of-the-box to create accurate digital twin environments for robotic manipulation tasks, like object rearrangement, assembly, or tool use (Fig. 1)? The appeal of single-view approaches extends beyond mere convenience. In robotics contexts, such as when entering a new room or opening a drawer, collecting multiple viewpoints often requires physical movement that delays decision-making and introduces additional failure modes through the need for viewpoint planning and collision avoidance under space constraints [29, 30]. Besides algorithmic simplicity, in scenarios such as books on a shelf or objects in cluttered drawers, gaining additional viewpoints is impossible without first interacting with and rearranging the scene [31, 32, 33].
The existing real-to-sim literature reveals a clear gap in this progression from multi-view to single-view approaches. Works that leverage dense viewpoints or scene scans [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] demonstrate promising result but require assumptions about scene coverage that limit their practical applicability. Conversely, works that attempt single-view approaches [34, 35, 36, 37] only produce domain randomised environments instead of exact digital twins with complete geometry, do not perform physics simulation, or, while mentioning potential robotics applications, lack comprehensive robot experiments. We describe these works and their limitations in greater detail in Section III-G.
@@
We are thus led to our central research question: can recent advancements in single-view mesh reconstruction models be effectively applied to create accurate digital twin environments for online, real-time, and high-throughput robotic manipulation? To address this question systematically, we formulate desiderata that are particular to robotics and simulation for manipulation, concerning the type of input data, computational resources, and overall coherence of the reconstructed scene. We empirically evaluate these desiderata by conducting experiments on two representative robotics datasets: the YCB-Video dataset [38] featuring standardised manipulation objects, and the Aria Digital Twin dataset [39] containing naturalistic household scenes. Our evaluation focuses on tabletop manipulation scenarios and assesses eleven state-of-the-art single-view 3D reconstruction models - those capable of inferring entire object geometries from a single viewpoint - across five key metrics: reconstruction accuracy, collision-free reconstruction, physical stability, occlusion handling, and computational efficiency.
@@
Our aim is to inform robotics practitioners and computer vision researchers about existing capability gaps, helping them select tools suited to their tasks while being aware of current limitations. Consequently, inspired by robotics requirements, we focus our evaluation on reconstruction models that:
@@
Work with single-view inputs. It is generally impractical to assume that an object can always be observed from multiple meaningful viewpoints. Beyond the practical constraints mentioned above, single-view operation is essential for real-time robotics workflows where viewpoint acquisition time directly impacts task performance.
@@
Produce complete meshes for individual objects. We are interested in models that produce complete object shapes, not partial geometry such as depth maps or visible surfaces only. Most robotics simulation engines operate on mesh representations [40, 41, 42]. While other types of 3D representation, such as point clouds or voxels, are supported in some cases, meshes remain the de-facto standard for rigid body simulation. We note, however, that alternative representations such as 3D Gaussian Splatting (3DGS) [43] are beginning to be explored for direct simulation [44], potentially offering new pathways for combining perception and dynamics. While promising, such approaches remain early-stage and have not yet been demonstrated in robotic manipulation settings. Exploring their suitability for real-time interaction and physical reasoning remains an important avenue for future research.
@@
Are trained to perform category-agnostic reconstruction. Historically, 3D datasets have been scarce and small [45, 46, 47]. As a result, most 3D reconstruction approaches focused on category-specific reconstruction, such as chairs or cars, to limit the task’s scope. However, the diversity of objects encountered in real-world robotics scenarios necessitates reconstruction models that can handle heterogeneous object classes without prior knowledge of scene contents. Recent developments in large-scale 3D datasets [48] and 2D data for for 3D reconstruction [49, 50] have enabled category-agnostic models that promise to reconstruct complete 3D geometry - including unobserved parts - across diverse object categories, making them particularly suitable for the varied environments encountered in household and industrial robotics.
Our systematic evaluation of works fitting these criteria complements existing surveys that attempt to exhaustively categorise 3D reconstruction models either generally [51, 52, 53, 54], or with a focus on specific 3D representations such as NeRFs [55], 3DGS [56, 57], or Meshes [58, 59]. We are not looking to provide an exhaustive overview of the 3D reconstruction model zoo but instead fundamentally address the question whether single-view reconstruction performance of existing models is sufficient to be used in robotics. While [57, 60] offer extensive overviews of 3D reconstruction in robotics, they mostly focus on scanning approaches that require substantial scene coverage. In contrast, we focus specifically on single-view reconstruction models and run experiments on real-world datasets to evaluate whether the reported performance on computer vision datasets transfers to realistic robotics domains. [61] have a somewhat similar motivation but focus on three text-to-3D generative models for mobile robotics, and investigate several optimisations to improve the latency vs reconstruction quality trade-off. We instead evaluate a broad and representative range of image-to-3D models for real-to-sim reconstruction.
In Section II, we delineate desiderata for single-view 3D reconstruction models that robotics domains necessitate. While then provide a brief taxonomy of existing 3D reconstruction approaches in Section III. Even though this work does not focus on multi-view approaches, such works are included for completeness and we refer the reader to relevant surveys, where appropriate. In Section IV, we systematically evaluate a selection of single-view 3D reconstruction models towards these desiderata to see to what extent existing models are readily applicable in the robotics domain. Finally, in Section V we summarise and discuss our findings and position them in the broader context of current research directions.
@@ II Desiderata
For robotic manipulation tasks such as grasping, placing, and rearranging objects on tabletops, existing single-view, category-agnostic 3D reconstruction models must fulfil several key requirements to ensure that the reconstructed models are not only visually accurate but also physically meaningful, allowing for realistic interactions within a simulation. Furthermore, deployment in the real world results in constraints regarding the available input data and available computational resources. In this section, we motivate several requirements that are key for digital twinning on robotic platforms in the wild.
@@
Robot manipulation is highly sensitive to accuracy in perceptual system components. For successful robot grasping and manipulation, surface estimates should be within 1mm of the true object surface [62, 63]. Some use-cases, such as assembly or part insertion, can even require sub-millimetre tolerances [64, 65, 66]. For our evaluation setting of household environments, we set the target accuracy to 2mm.
These stringent accuracy requirements, however, may be challenging to meet with current 3D reconstruction models due to a potential domain gap in training data. Real-world robots are often equipped with RGB-D cameras which offer both RGB and depth perception at relatively low cost [67]. Yet, these cameras typically do not match the resolution and colour profile of images commonly found in 3D reconstruction training datasets [47, 48]. The difference is especially noticeable when an object occupies only a small part of the image plane, as is common when observing an entire scene. Consequently, the 3D reconstruction training data distributions [68, 69, 70] differ substantially from the distribution of inputs encountered during deployment in the real world, potentially resulting in less accurate reconstructions and artifacts [15]. While some approaches like [71] co-train on both real-world images and synthetic data to improve reconstruction accuracy, it remains an open question whether this approach can yield satisfactory results for the precision demands of robotics applications.
@@
Physics simulators typically resolve collisions through penalty- or constraint-based methods that modify the current force acting on the object, its velocity, or its position [40, 41, 42]. Many 3D reconstruction approaches process objects individually which can result in collisions when placing these reconstructions back into the scene [68, 69, 70]. This is a direct result of insufficient supervision signal from the immediate vicinity of the reconstructed objects. Scenes that collide already during static scene reconstruction break down when performing physics steps in the simulator due to the forces and constraints that act upon colliding objects. In principle, one could selectively deactivate collisions between objects to prevent this but that could lead to further complications in downstream simulation and cannot resolve situations where collisions are necessary to preserve the integrity of a scene, such as objects resting on supporting surfaces.
Desideratum 2: No object should violate space occupied by other objects or parts of the environment. This is ensures physical stability and accurate interaction simulation.
@@
This desideratum is based on the assumption that observed scenes are static as long as nothing interacts with them. We can consequently assume that all objects are in a stable, resting state and do not move when advancing simulated physics [72, 73, 74, 75]. Object pose and an object’s surroundings are necessary conditioning factors as otherwise the reconstruction process could ignore object and scene parts that are critical for object stability in the scene. Especially occluded object parts and parts with little surface area tend to be ignored during reconstruction, and such inaccuracies can lead to physical instability [72].
Desideratum 3: The reconstructed meshes should offer physically stable poses within 5 tilt of their scene pose, meaning that all objects are in a stable resting state and the integrity of the scene is maintained when advancing simulated physics.
@@
Reconstruction models must accurately resolve partial object occlusions that are ubiquitous in real-world environments [33, 76]. Without proper occlusion handling, digital twins would contain incomplete objects, leading to incorrect simulation physics and failed manipulation attempts. Some 3D reconstruction models run a separate occlusion completion model in their pipeline which can lead to computational overhead and accumulation of reconstruction error [35]. Other works assume that objects tend to be symmetrical or follow other manual shape priors [77, 78]. Object stability, as described in the previous desideratum, can give valuable clues about the structure of occluded parts.
Desideratum 4: For common objects, such as the household objects represented in our evaluation datasets, the reconstructed meshes should resolve partial occlusions with Chamfer distances no more than 10% higher than those of visible regions when compared to ground truth, ensuring accurate simulation of partially-occluded objects.
@@
Many 3D reconstruction models that rely on iteratively optimising a shape representation take several minutes or even hours to reconstruct objects or scenes [70, 79, 72, 75]. As this reconstruction pipeline needs to be run every time a new object is observed, it represents a significant computational cost when deploying (mobile) robots in online environments. Other approaches are able to produce shape reconstructions within few seconds by relying on a single forward pass [27, 80, 68, 26, 81]. Naturally, latency requirements differ considerably with the type of task that is to be completed. For real-time systems operating at multiple-Hertz frequency [82, 83], however, these models are likely too slow, especially since compute time accumulates if multiple objects are to be reconstructed. We define a target latency of 2s for reconstructing an entire scene consisting of multiple objects, because the evaluation data used in this paper concerns less time-critical applications, such as household environments. While one could argue that pre-computing reconstructions and retrieving them from a database when needed might address computational concerns, this approach is generally impractical in online environments like households where object positions, orientations, and even the objects themselves change frequently. Another perspective is given by a possible accuracy/efficiency trade-off, as some computationally more expensive models produce lower-error reconstructions. This would indicate that the practitioner has to make a judgement call as to what reconstruction error and computational cost can be afforded.
Additionally, robots generally have limited on-board compute due to power and payload restrictions [84]. At the same time, many 3D reconstruction models utilise computationally expensive operations, such as diffusion [27, 69] or iterative optimisation [70, 49, 79], rendering them unable to be run on-device without significantly increased latency. Running reconstruction models on dedicated servers and streaming the results to the robotic platform is often not feasible due to latency issues, network dependence, and potential security or privacy concerns [85].
Desideratum 5: The reconstruction pipeline should be able to produce a complete scene reconstruction within 2s, even when multiple objects are present, ensuring efficient reconstruction for online, real-world environments.
@@
In this section, we provide a focused literature overview of 3D reconstruction approaches to motivate our approach and position our findings. Given the rapidly evolving nature of this research area with new methods appearing frequently, we cannot comprehensively cover all existing approaches. Instead, we focus on representative methods that illustrate key capabilities and limitations relevant to robotic manipulation requirements. This framework highlights how current approaches align with or fall short of the practical needs of robotic systems operating in real-world environments. Specifically, we look at different 3D representations that are used for reconstruction (Section III-A), various input data requirements (Section III-B), different prediction paradigms (Section III-C), different representation abstraction levels (Section III-D), the consideration of physical properties and parameters during reconstruction (Section III-E), training data properties (Section III-F), and how recent reconstruction approaches have been leveraged in a real-to-sim fashion in robotics (Section III-G). Table I gives an overview of a selection of existing reconstruction models and their properties.